a rectangular area is 5 m longer as it is wide. one side of a square swimming pool will match exactly the width of the area . after the pool is built the area left will be 30 m squared. how wide is the area
Answer by ankor@dixie-net.com(22740)

a rectangular area is 5 m longer as it is wide.
one side of a square swimming pool will match exactly the width of the area .
after the pool is built the area left will be 30 m squared.
how wide is the area?
:
Let x = the width of the pool and the side of the square pool
:
Write an area equation
Rect area - pool area = 30 sq/m
x(x+5) - x^2 = 30
x^2 + 5x - x^2 = 30
x^2 - x^2 + 5x = 30
5x = 30
x = 30/5
x = 6 meters is the width
;
:
Check this, find the rect area and the pool area (rect length 6+5)
11 * 5 = 55 sq
5 * 5 = 25
--------------
left over 30 sq/ft